OpenStudy (anonymous):

A bus accelerates at -0.5 m/s2 for 12.5 seconds. If the bus had an initial speed of 31 m/s, what is its new speed? A. 37.25 m/s B. 6.25 m/s C. 24.75 m/s D. 12.25 m/s

OpenStudy (cybershadow):

Given a = -0.5 m/s^2 t = 12.5 seconds u = 31 m/s v = ut + 1/2 at^2 v =31 * 12.5 + 1/2 * (-0.5) * (12.5)^2 v = 348.43 m/s Weird! , Are u sure ur options are correct?

OpenStudy (cybershadow):

or u have mistyped something in question?

OpenStudy (cybershadow):

oh nevermind

OpenStudy (cybershadow):

hahah wat A Huge Mistake , BRAINDEAD

OpenStudy (cybershadow):

v = u + at

OpenStudy (cybershadow):

so v = 31 * 12.5 + (-0.5) * 12.5 = 24.75 m/s

OpenStudy (isaiah.feynman):